-VP use loplace transform to solve y-ayt y = f(t), 9 (6)=1, 96)=1 f(t) o tas It- 5t25 you may use partial de comp but must
Answer #1

y-2y +y = f(t) solution: t<5 PCt). Poi yo=1 16-5;t 2,5 y co=1 LCyct)) SyCs) - syco) - YCC) LCyct)) - SY() - Yo L (y(t)) =yos) -5S + (6-) $225+1) (2_25+) -55 t sal (saya 32 (soja -55 = e 와 s-1 (s-ija S A + B + s. S-1 5 5.2 & (39-25+1) §§ (G-172 .(Therefore, C=-2B C=-2 B=1 B+C+D = O 1-2+D Del 2 t (599+) sisija (6-D2 S- Therefore, -58 Y()-é - + 2 5-1 ( S-1 G-12 -55 -SS YG